Apr 10, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and SeeVogh session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting May 12-14
  3. Action Items
  4. FCAL
  5. BCAL
  6. Reconstruction and Simulation
  7. TDR update
  8. Calibration
  9. Any other business


Attendees: Elton, John, Will, Yi, David, Mark (JLab); Shaun, Andrei, Zisis (UofR); Matt (IU)

  1. Announcements
    • CDC has been inserted into bore of the magnet. Trial cabling is ongoing, which makes access to the upstream BCAL more difficult. It will be backed out for surveys next week.
  2. Collaboration Meeting May 12-14
  3. Action Items
  4. FCAL (John)
    1. Box of bases has been returned to IU for diagnosis and repair
    2. Detector nearing 100% ready, minus 2-3 channels which are misbehaving.
    3. FADC issues
      • New CODA has been provided by Serguei Furletov and is working well. Full checkout installation allows for more control.
      • Many entries with missing data were found to have late pulses by 40 - 50 samples. Timing is being investigated and confusing experts that don't see how this can happen for individual channels (not entire modules)
      • Some FADC channels are still stuck at fixed values
      • High pedestals in some channels have disappeared since new CODA has been installed, although FADC firmware has not been updated.
    4. Matt: Can one tell if failures are simply due to infant mortality or slow continuous failure rate? John: Cannot tell. They have run the system over night with no failures, but found some failures after quiet weekend without taking data. Failures could be due to turning the detector on/off.
    5. David: Asked if BCAL has seen the frozen FADC values. Elton: No, but we have taken very little data.
  5. BCAL
    1. Detector checkout
      • Elton and Will checked "dead" channels from the two north crates that were read out.
      • Three channels had loose Lemo connectors
      • Two channels failed to detect the LED signal because it was too small. (Nominal setting for LED bias has been 6V). The signals appeared after increasing the voltage to 7 V.
    2. [Checking interlocks]
    3. [Restoring HV from file]
      • Zisis: Could create standard files for specific overvoltages and temperatures.
      • David: Asked about low voltage modules loosing settings and restoring to damaging settings. Elton: This is being investigated, but at least on the BCAL the low voltage modules are the MPOD modules that retain their settings under power loss.
      • Zisis: We should check the max voltage settings for the MPPCs. Elton will do this.
      • Yi: On the note that some RTDs had transient temperature readings for several seconds. This is probably not physical. Elton: Could be that there is still a connection between voltages and the RTDs that shows up as a transient in the temperature readings.
    4. Slow Controls
      • The LV and HV GUIs for the BCAL are being used routinely now. Incremental improvements are making them easier to use.
      • Orlando has implemented a GUI for the LED pulser system and is also working well at this time.
    5. DAQ readout
      • New implementation of DAQ from Furletov is almost ready. (Works well for the FCAL
      • The DAQ is now reading the four FADC crates on the north side
      • Readout list in crate 31 had insufficient memory allocation
      • Will take new data set after lunch
      • Zisis: Once most of these issues have been cleared up, Andrei can import longer files to Regina for analysis.
    6. ESB Cosmics (Shaun)
      • Elton supplied the last four files from the ESB cosmic tests. Shaun will analyze to complete his master's thesis. He will report on this next time.
  6. Reconstruction and Simulation
  7. TDR update
    • Latest edits by Zisis were missing files, but these have been checked in now
    • Elton will look at the draft next.
    • We need to include some text regarding expected life time of MPPCs due to irradiation. Yi has uploaded his talk at the S&T review to the portal, which has some of the details.
    • FCAL draft is on Matt's "to do list"
  8. Calibration
  9. Any other business